### v1

@@ II.6 Parameter sets
The steps from the previous subsections are simulated in NetSquid to find the error probability (i.e., the fraction of failed test rounds) for a given setup. The simulation takes a set of hardware parameters as input. We divide these hardware parameters into two sets: one set that we do not vary during optimisation (parameter set 1), and one set that we do vary during optimisation (parameter set 2). The first set contains parameters that are inherent to the setup (fibre length), parameters that are established optical components with little room for improvement (fibre loss, crosstalk in beamsplitters, waveplate error probability, detector dark-count rate) and the duration of operations at the ion trap. The last is not optimised over as these timescales are already very small compared to the time required to perform RSP (which is in the order of 20 ms per qubit, two orders of magnitude larger than the timescale of performing gates and readout). This first set is presented in Table 1.
@@
Channel length
50 km
Photon loss probability in fibre
0.2 dB/km
Waveplate error probability (fast axis tilt/retardation deviation)
0.001
Dark count probability photondetectors
0.03%
Crosstalk in polarising beam splitter
0.0001
Qubit rotation duration
12 $\mu$s [44]
Entangling gate duration
107 $\mu$s [44]
Ion initialisation duration
300 ns [45]
Photon emission duration
300 ns [45]
Ion qubit readout duration
100 $\mu$s
Table 1: Parameter set 1: Hardware parameters used as simulation input that are not varied over.
The second set of parameters are the optimisation parameters. These are the parameters we vary over in the minimisation methods explained in Section IV.3. In Table 2 we list these parameters as well as what we consider the ‘baseline’. The baseline is current state of the art in long-distance trapped-ion experiments and is used in the minimisation methods to find the difference between what is currently achievable and what is needed to run the protocol at metropolitan distances, as will be explained in Section IV.3. We choose this baseline instead of values from current record experiments (such as the long coherence time of [46], the high gate fidelities of [34, 32] or the high entanglement fidelity of [47]) because we believe it is more realistic for long-distance experiments. We are actively looking ahead to what we need to achieve rVBQC in a metropolitan scale (i.e., over a 50 km distance), instead of focusing on in-lab experiments, as rVBQC is already feasible in a lab setting [22].
@@ 12 $\mu$s [44]
(= emit $\times$
freq. conversion)
@@
(= 0.53[48] $\times$
0.25[49])
Single-qubit gate fidelity
0.99 [50]
Entangling gate fidelity
0.95 [50]
Emission fidelity
0.974 [51]
Coherence time [ms]
62 [44]
Table 2: Parameter set 2: Baseline of long-distance trapped-ion experiments forming state of the art. We vary over these parameters to find the absolute minimal hardware requirements and minimal improvements as outlined in Section IV.3. Server efficiency here refers to the total efficiency of preparing an ion-qubit, emitting an ion-entangled photon, coupling to the fibre (combined giving the number for ‘emit’ in the table) and converting its frequency to the telecom C band at 1550 nm (‘freq. conversion’ in the table). The emission fidelity refers to the fidelity of the ion-photon entangled pair when the photon is emitted.
Figure 4: Absolute minimum a given parameter can be while still staying below the required maximum error probability of 25% assuming that the only other imperfection in the system is photon loss in fibre (orange). These minima are compared to the state of the art as given in Table 2 (blue/purple, hatched). The dashed black line indicates a change in scaling between the fidelities and the coherence time. Below the solid black line there is a comparison between the overall error with all imperfections in the state of the art (yellow) and the threshold error probability of 25% (green hatched); despite each parameter’s baseline value meeting the minimal requirements, together they are insufficient to run the rVBQC protocol.
@@ II.7 Error tolerance and requirements
In abstract cryptography (i.e., the theory behind the security proof of the rVBQC protocol [20]), security is defined as the indistinguishability between the real-world implementation of the protocol and the ideal, noiseless resource [52]. When the security of a protocol holds over sequential or parallel repetitions with other protocols, it is said to be composably secure. The rVBQC protocol we study (A) is $\epsilon$-composably-secure (as defined in [20]) with $\epsilon$ exponentially small in the number of rounds if the fraction of failed test rounds over total number of test rounds is bounded by 25% [20]. Details for this are provided Section IV.3.
We can use the error tolerance to find two sets of requirements: absolute minimal hardware requirements and minimal improvements. To find the absolute minimal hardware requirements, we start by setting all hardware parameters to perfect except for photon loss in fibre. That is, for parameter set 1, we set the waveplate errors, darkcount probability and beam splitter crosstalk to 0, the loss in fibre is kept as presented in Table 1 at 0.2 dB/km. For parameter set 2, we set the server efficiency and all fidelities to 1 and we remove the memory decoherence noise model to simulate effectively infinite coherence time. Then one parameter from set 2 is made progressively worse, until the simulated error probability rises above 25%. The last value of the parameter before this happens is taken as the absolute minimal requirement.
For the minimal improvements, we aim to find the least costly improvements needed over the baseline parameters to get the error probability below 25%. The cost of a hardware parameter is given in terms of an improvement factor. This improvement factor quantifies the difference between a given value and the baseline, being 1 for no improvement, and tending to infinity for a parameter tending to perfect (e.g., infinite coherence time, fidelity of 1). More information on the cost function and the improvement factor can be found in Section IV.3. The minimal improvements are found by combining the error tolerance with minimising the improvement factors into a single objective optimisation problem and solving it using a genetic algorithm.
@@ III Results and discussion
Figure 5: (a): Minimally improved parameter set required to perform a 5-qubit linear-graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km with a 25% error probability threshold. This set minimises the cost function (11), meaning it is closest to the state-of-the-art baseline of those that meet the requirement. A visual representation of these parameters in terms of an improvement factor is given in the bottom figure. (b): Directions along which hardware must be improved to perform a 5-qubit linear graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km. The further away the line is from the centre towards a given parameter, the larger the improvement that parameter requires. The improvements are given in terms of an improvement factor k, which tends to infinity as a parameter tends to its perfect value and is 1 for no improvement. More information on this can be found in Section IV.3.
@@
We now find requirements for enabling rVBQC with a measurement-only client and a trapped-ion server separated by 50 km of optical fibre, for five-qubit linear graphs. The data generated for these results is available at [53], the code used to generate the data and plots is available at [54].
The absolute minimal requirements found, compared to the baseline of Table 2, are shown in Figure 4. More information on how the values of Figure 4 are extracted from the simulation outcomes and on how their confidence intervals are determined is given in Section IV.3.
From Figure 4 we see that state-of-the-art ion traps satisfy absolute minimal requirements on a per-parameter basis as all blue bars, representing the baseline, are larger than the absolute minimal requirements in orange. However, in Figure 4 we also show the comparison between the full set of baseline parameters as ‘overall error’ in red (i.e., all imperfections combined), with the threshold in green. This shows us that all current imperfections combined make it impossible to perform the blind computation securely over 50 km using existing technology, as all baseline parameters combined lead to an error probability of 50%, which is twice the requirement of 25%.
@@
Note that there is no absolute minimal requirement on the server efficiency, as there is no threshold for rate, only for error probability. If all other parameters are perfect, any server inefficiency combined with fibre attenuation will still lead to a 0% error probability, as the storage of qubits is perfect. (This is not the case for the minimal improvements set, as other parameters will be imperfect. Having imperfect memory efficiency then leads to qubits residing in imperfect memory for longer, thereby suffering more decoherence.) It will only lead to lower rate, not considered here. The difference between the minimal requirement and the state of the art appears particularly large for the coherence time, however, we note that the comparison is skewed as all other parameters are on a scale of zero to one, which is not the case for coherence time. Because of this we chose to separate the coherence time from other parameters with a dashed line, to indicate a different scaling.
@@
We find the minimal improvements, i.e., the set that is closest to the baseline in Table 2 satisfying the threshold, by setting up and minimising a cost function as defined in (11). More details on the cost function and how the minimisation is performed can be found in Section IV.3. The set of hardware parameters that minimises this cost function are given in Figure 5 (a) and visualised in Figure 5 (b). The plot in Figure 5 (b) shows the improvement factor for each parameter in Figure 5 (a). The improvement factor tends to infinity as a parameter tends to its perfect value and is 1 when no improvement is made compared to the baseline. The further away the line is from the centre towards a given parameter, the larger improvement that parameter requires. More information about this is also given in Section IV.3.
From this we can see that comparatively little improvement is needed in server efficiency and emission fidelity. The server efficiency for the ion trap considered here is enhanced due to the use of a cavity [48]. Optimising this further is less important in reaching the error probability threshold than increasing the coherence time because we are in a regime where the time most qubits spend in memory is limited by the cutoff, thereby increasing the efficiency will lead to a higher rate (not considered here) rather than a lower error probability.
@@
More improvements are needed for the entangling and single-qubit gate fidelities, and coherence time. As can be seen from Figure 3, many single-qubit gates are executed in order to perform the $CZ$ gate used for creating the graph state, which makes any imperfections in the execution of such single-qubit gates more impactful. We do note that the number of single-qubit gates may be reduced by optimising the graph state creation and perhaps absorbing some of the gates into the measurement bases, which would reduce the overall errors and hence lessen the need for improvement on this parameter. This is not the case for the entangling-gate fidelity.
@@ IV Methods
Here, we discuss some details of how the trapped-ion server and client apparatus are modelled in our NetSquid simulation available at [54]. We also outline the minimisation method used to determine the requirements mentioned in the previous section.
@@ IV.1 Trapped-ion modelling
We model the trapped-ion server using the NetSquid-based library NetSquid-TrappedIons [25]. This library has been first used in [27] and is here introduced in more detail.
We model the decoherence of trapped-ion qubits over time using the collective Gaussian dephasing process described in [55], which can be rewritten as [27]
@@
The emission of entangled photons from an ion trap is modelled in NetSquid-TrappedIons as the creation of a photon that has a polarisation degree of freedom that is maximally entangled with the state of the ion used in the emission, |$\ket{\psi}=(\ket{0H}+\ket{1V})/\sqrt{2}$, followed by the application of a single-qubit depolarising channel on the photon’s polarisation. This results in a Werner state of the form $\frac{3}{4}F\ket{\psi}\bra{\psi}+\frac{4F-1}{3}\frac{\mathbb{1}}{4}$, where $F$ is the fidelity of the state to the perfect state $|\psi\rangle$.
@@ IV.2 Client apparatus modelling
The client is modelled as an optical setup where an incoming ion-entangled photonic qubit is transmitted through a half waveplate (HWP) followed by two quarter waveplates (QWPs) to rotate the polarisation state. The photon is then guided to a polarising beam splitter and detected with single-photon detectors.
In the simulation, the effect of the waveplates is given by the multiplication of the waveplate Jones matrices (IV.2) with the state vector [56]. The relative phase retardation induced between the fast axis and the slow axis is $\delta=\pi/2$ for a quarter waveplate and $\delta=\pi$ for a half waveplate. The fast axes waveplates also have an angle $x$ with respect to the x-axis, determining the specific rotation that is implemented. However, errors in the setup can influence the retardation, giving a retardation deviation of $\Delta\delta$. It is also possible to have a deviation in the angle $x$ leading to $\Delta x$. With this, we can write the Jones matrices in general form [57] to include the errors as
@@
where $\delta^{\prime}=\delta+\Delta\delta$ (with $\delta=\pi/2$ for a QWP, and $\delta=\pi$ for a HWP) and $x^{\prime}=x+\Delta x$. Together, $\Delta\delta$ and $\Delta x$ lead to the waveplate error probability as given in Table 1. The waveplates are implemented as a custom operation in NetSquid according to the Jones matrix.
[30] gives an overview of general qubit rotations in terms of these fast axis settings. We can set the fast axes to correspond to a measurement in the |±$|\pm_{\theta}\rangle$-basis as
$x_{1}=0;$
$x_{2}=\theta/2;$
$x_{3}=\theta/4-3\pi/4.$
@@ IV.3 Minimisation methods
In our analysis, we use one target metric. This target metric is the error tolerance of 25% that has been introduced before. This is a bound on the fraction of test rounds that are allowed to fail while still being $\epsilon$-composably-secure. In [20] the fraction of failed test rounds $w$ over the total number of test rounds $t$ is bound by
@@
Here, $k$ is the principal colouring of the computation graph (which is the smallest number of ‘colours’ or labels one can give to a the nodes in a graph to where no two neighbouring nodes have the same colour, see for example [58]) and $p$ is the inherent error probability of the bounded-error quantum computation. If the error probability is above this fraction, the client aborts the protocol due to security issues. Assuming a computation for which $p=0$ we require the error probability to be below $1/2k$. The one dimensional graph for single-qubit rotations used in this paper is two-colourable, which gives a maximum error tolerance of 25%.
@@
We look for absolute minimal hardware requirements by setting all-but-one parameters to perfect aside from fibre attenuation. For the coherence time, perfect means removing the collective dephasing noise model from the ions in the trap. We then sweep over the non-perfect parameter to find where the error probability due to this imperfection crosses the threshold.
@@
To find the crossing point, we do an initial global search with a small number of test rounds per point (to limit computation time), to find the approximate regime of where the error probability would pass the threshold. Once the regime is located, the search is focused by taking the closest point above and below the threshold, halving the distances to their mean and running the simulation again for these points with a larger number of rounds. This process is repeated until the error-probability confidence intervals of the points crosses the threshold of 25% (this is similar to the bisection method used in root finding). The confidence intervals are determined by Hoeffding’s bound [59] as $\sqrt{\log{(2/0.05)}/2t}$ with $t$ the number of test rounds. The focused search is executed with 70000 points, in order to have a confidence level of 95% in an interval $\pm$ 0.005 for the error probability. The minimal requirements given in Figure 4 are then extracted from the closest points ($(x_{1},y_{1}),(x_{2},y_{2})$) by a linear interpolation at the threshold ($y=0.25$) as $x=(y-y_{1})*(x_{2}-x_{1})/(y_{2}-y_{1})$. The error in these estimates is found by applying the same interpolation to the edges of the error probability confidence interval.
@@
Next, we find the set of minimal improvements. That is, from a given baseline (Table 2), what parameters allow for fulfilling the constraint on the error probability with the least improvement. In order to quantify the cost of improving a parameter by a certain amount, we define hardware costs $H_{c}$ and a cost function $C$, which combines the hardware costs with our constraint on the error probability to give a single-objective minimisation problem as done in [26, 60, 61, 27]. We then employ a genetic algorithm using a workflow manager called YOTSE [62] to find the set of hardware parameters that minimises the cost function. This optimisation was run on SURF’s high- performance-computing cluster Snellius (Platinum 8360Y CPU @ 2.4GHz, maximum of 480 GB RAM) and on a workstation featuring an Intel Xeon Gold 6230 CPU @ 2.10GHz and 188 GB of DDR4 RAM, taking around 100 hours per optimisation run of 20 generations.
@@
This is equivalent to summing over the improvement factor of each parameter.
The probabilities of no-imperfection are defined for the optimisation parameters as in Table 3.
@@
Server efficiency $\eta$
$\eta$
Coherence time $T_{c}$
$e^{-t^{2}/T_{c}^{2}}$
(Gate and entangled state) fidelity $F$
$\frac{1}{3}(4F-1)$
@@
For derivations and more explanation of the probabilities of no-imperfection, see Supplementary Note 6 of [27]. Note that the $t$ in the probability of no-imperfection of the coherence time, indicating the timescale over which the decoherence happens, does not influence the hardware cost, as $\ln\left(p_{NI}(b_{i})\right)/\ln\left(p_{NI}(x_{i})\right)$ = $\ln\left(e^{-t^{2}/T_{c}^{2}}\right)/\ln\left(\sqrt[k]{e^{-t^{2}/T_{c}^{2}}}\right)$ = $\left(-t^{2}/T_{c}^{2}\right)/\left(-t^{2}/kT_{c}^{2}\right)=k$.
@@
We now combine our requirement of having an error probability below 25% with the hardware cost to find the total cost of a set of parameters. We want the cost assigned to a set of parameters to be very high when the constraint is not met, and to be lower the closer parameter sets are to the baseline assuming that the constraint is met. A function to capture this behaviour, similar to what is used in [26, 27, 60, 61], is
$\begin{split}C=&w_{1}(1+(w/t-1/2k)^{2})\Theta(w/t-1/2k)\\ &+w_{2}Hc(x_{1},...,x_{N}),\end{split}$ (11)
@@
where $\Theta(x)$ is the Heaviside step function and $w_{1}$ and $w_{2}$ are the weights of the objectives. We choose $w_{1}\gg w_{2}$ in order for the function to reflect that it is much more important to satisfy the requirement than to minimise hardware requirements, i.e., we do not care about the hardware cost as long as the requirement is not met.

### v2

@@ II.6 Parameter sets
The steps from the previous subsections are simulated in NetSquid to find the error probability (i.e., the fraction of failed test rounds) for a given setup. The simulation takes a set of hardware parameters as input. We divide these hardware parameters into two sets: one set that we do not vary during optimisation (parameter set 1), and one set that we do vary during optimisation (parameter set 2). The first set contains parameters that are inherent to the setup (fibre length), parameters that are established optical components with little room for improvement (fibre loss, crosstalk in beamsplitters, waveplate error probability, detector dark-count rate) and the duration of operations at the ion trap. The last is not optimised over as these timescales are already very small compared to the time required to perform RSP (which is in the order of 20 ms per qubit, two orders of magnitude larger than the timescale of performing gates and readout). This first set is presented in table 1. We do not take into account errors due to polarisation drift in fibre. This effect is assumed to be very small due to corrections with techniques such as the fully automated stabilisation using reference pulses described in reference [53].
@@
Channel length 50 km
Photon loss probability in fibre 0.2 dB/km
Waveplate error probability (fast axis tilt/retardation deviation) 0.001
Dark count probability photondetectors 0.02%∗∗
Crosstalk in polarising beam splitter 0.0001
Qubit rotation duration 12 $\mu$s [54]
Entangling gate duration 107 $\mu$s [54]
Ion initialisation duration 300 ns [55]
Photon emission duration 300 ns [55]
Ion qubit readout duration 100 $\mu$s
Table 1: Parameter set 1: Hardware parameters used as simulation input that are not varied over. The waveplate error probability is discussed in section IV.2. ∗∗ The dark count probability is given as $1-e^{-R_{\text{dc}}\tau}$ with $R_{\text{dc}}\approx 1500$ Hz the dark count rate for SPDs such as the SPDMA Si Avalanche Photodetector by Thorlabs and $\tau=12.5$ ns the detection time window such as in the supplementary material of Reference [35].
The second set of parameters are the optimisation parameters. These are the parameters we vary in the minimisation methods explained in section IV.3. In Table 2 we list these parameters as well as what we consider the ‘baseline’. The baseline reflects the current state of the art in long-distance trapped-ion experiments and is used in the minimisation methods to find the difference between what is currently achievable and what is needed to run the protocol at metropolitan distances, as will be explained in section IV.3. We choose this baseline instead of individual values from current record experiments (such as the long coherence time of reference [56], the high gate fidelities of references [47, 45] or the high entanglement fidelity of reference [57]) because we believe it is more realistic for long-distance experiments. We are actively looking ahead to what we need to achieve rVBQC at a metropolitan scale (i.e., over a 50 km distance), instead of focusing on in-lab experiments, as rVBQC is already feasible in a lab setting [35].
@@
(= emit $\times$ freq. conversion)
@@
(= 0.53[58] $\times$ 0.25[59])
Single-qubit gate fidelity 0.99 [60]
Entangling gate fidelity 0.95 [60]
Emission fidelity 0.974 [61]
Coherence time [ms] 62 [54]
Table 2: Parameter set 2: Baseline of long-distance trapped-ion experiments consistent with the state of the art. We vary over these parameters to find hardware requirements to perform rVBQC. Server efficiency here refers to the total efficiency of preparing an ion-qubit, emitting an ion-entangled photon, coupling to the fibre (combined as the number for ‘emit’ in the table) and converting its frequency to the telecom C band at 1550 nm (‘freq. conversion’ in the table). The emission fidelity refers to the fidelity of the ion-photon entangled pair when the photon is emitted.
Figure 4: Absolute minimum value a given parameter can have that is consistent with the required maximum error probability of 25%, assuming that the only other imperfection in the system is photon loss in fibre (orange). These minima are compared to the state of the art as given in Table 2 (blue/purple, hatched). The dashed black line indicates a change in scaling between the fidelities and the coherence time. Below the solid black line, the overall error with all imperfections in the state of the art (yellow) is compared with the threshold error probability of 25% (green hatched); despite each parameter’s baseline value meeting the minimal requirements, together the baseline values are insufficient to run the rVBQC protocol.
@@ III Results and discussion
Figure 5: (a): Minimally improved parameter set required to perform a 5-qubit linear-graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km with a 25% error probability threshold. This set minimises the cost function (10), meaning it is the closest to the state-of-the-art baseline of the sets that meet the requirement. (b): Directions along which hardware must be improved to implement a 5-qubit linear graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km. The further away the line is from the centre towards a given parameter, the larger the improvement that parameter requires. The improvements are given in terms of an improvement factor $k$, which tends to infinity as a parameter tends to its perfect value and is 1 for no improvement.
In abstract cryptography (i.e., the theory behind the security proof of the rVBQC protocol [32]), security is defined as the indistinguishability between the real-world implementation of the protocol and the ideal, noiseless resource [62]. When the security of a protocol holds over sequential or parallel repetitions with other protocols, it is said to be composably secure. The rVBQC protocol we study (A) is $\epsilon$-composably-secure (as defined in [32]) with $\epsilon$ exponentially small in the number of rounds if the fraction of failed test rounds over total number of test rounds is bounded by 25% [32]. Details for this are provided in section IV.3. The protocol includes classical error correction in the form of a repetition code, but includes no quantum error correction. Note that the 25% error constraint is only present for rVBQC, in non-verifiable BQC the protocol can in principle be performed at any error probability but without any guarantee of the correctness of the outcome.
We can use the error tolerance to find two sets of requirements: absolute minimal hardware requirements and minimal improvements. To find the absolute minimal hardware requirements, we start by setting all hardware parameters to perfect except for photon loss in fibre. That is, for parameter set 1, we set the waveplate errors, dark-count probability and beam splitter crosstalk to 0. The loss in fibre is kept as presented in Table 1 at 0.2 dB/km. For parameter set 2, we set the server efficiency and all fidelities to 1, and we remove the memory decoherence noise model to simulate effectively infinite coherence time. Then one parameter at a time from set 2 is made progressively worse, until the simulated error probability rises above 25%. The last value of the parameter before this happens is taken as the absolute minimal requirement.
For the minimal improvements, we aim to find the least costly improvements needed over the baseline parameters to get the error probability below 25%. The cost of a hardware parameter is given in terms of an improvement factor. This improvement factor quantifies the difference between a given value and the baseline, being 1 for no improvement, and tending to infinity for a parameter tending to perfect (e.g., infinite coherence time, fidelity of 1). More information on the cost function and the improvement factor can be found in section IV.3. The minimal improvements are found by combining the error tolerance with the improvement factors into a single objective optimisation problem and solving it using a genetic algorithm.
@@
We find requirements for enabling rVBQC with a measurement-only client and a trapped-ion server separated by 50 km of optical fibre, for five-qubit linear graphs.
@@
The absolute minimal requirements found, compared to the baseline of Table 2, are shown in figure 4. We see that state-of-the-art ion traps satisfy absolute minimal requirements on a per-parameter basis. However, we also calculate the error probability with the full baseline of parameter set 1. This shows us that all current imperfections combined make it impossible to perform the blind computation securely over 50 km using existing technology, as the corresponding error probability of 51% is twice the requirement of 25%. Note that there is no absolute minimal requirement on the server efficiency, as there is no threshold for rate, only for error probability. If all other parameters are perfect, any server inefficiency combined with fibre attenuation will still lead to a 0% error probability, as the storage of qubits is perfect. (This is not the case for the minimal improvements set, as other parameters will be imperfect. Having imperfect server efficiency then leads to qubits residing in imperfect memory for longer, thereby suffering more decoherence.) It will only lead to lower rate, not considered here. The difference between the minimal requirement and the state of the art appears particularly large for the coherence time. However, we note that the visualisation is skewed as all other parameters are on a scale of zero to one, which is not the case for coherence time. Because of this we chose to separate the coherence time from other parameters with a dashed line, to indicate a different scaling.
@@
We find the minimal improvements, i.e., the set that is closest to the baseline in table 2 satisfying the threshold. The set of hardware parameters that minimises the cost function are given in figure 5 (a) and visualised in figure 5 (b). The plot in figure 5 (b) shows the improvement factor for each parameter in figure 5 (a). The further away the line is from the centre towards a given parameter, the larger improvement that parameter requires. From this we can see that comparatively little improvement is needed in terms of improvement factor for server efficiency and emission fidelity. The server efficiency for the ion trap considered here is enhanced due to the use of a cavity [58].
@@
More improvements are needed for the entangling and single-qubit gate fidelities and for the coherence time. As can be seen from figure 3, many single-qubit gates are executed in order to perform the $CZ$ gate used for creating the graph state, which makes any imperfections in the execution of such single-qubit gates more impactful. We do note that the number of single-qubit gates may be reduced by optimising the graph state creation and perhaps absorbing some of the gates into the measurement bases, which would reduce the overall errors and hence lessen the need for improvement on this parameter. This is not the case for the entangling-gate fidelity.
Notably, a lower improvement factor does not always means a value is easier to obtain, this lies in the problem of finding a meaningful cost function. In this case, most values have been obtained in separate optimised experiments, but this is not true for the server efficiency. The baseline used for this optimisation is already optimised for server efficiency as it comes from communication experiments, such that further improvement might be more challenging. Therefore, while the improvement factor might help in visualisation, the true values given in table of figure 5(a) might give a clearer idea of what improvements are needed.
Genetic algorithms do not guarantee to find a global minimum, instead several local minima were found, of which the one presented here was the lowest. Other solutions are roughly equivalent but might give slightly more importance on improving one parameter over the other. An alternative solution is given in B for comparison, and other datasets can be found in [63].
@@ IV Methods
Here, we discuss some details of how the trapped-ion server and client apparatus are modelled in our NetSquid simulation. We also outline the minimisation method used to determine the requirements identified in the previous section.
Though in principle the optimisation can be executed using a different simulator, the choice for using NetSquid as opposed to other quantum network simulators is threefold. First, it is well suited for this type of modelling, as it is a dedicated quantum network simulator that simulates the hardware layer on an appropriate scale. Second, an ecosystem of open-source user-contributed libraries has developed around NetSquid, providing us with useful tools and examples [64, 65]. We have been able to contribute back to this ecosystem by integrating our own library NetSquid-TrappedIons [38]. Lastly, some of the authors of this work having prior experience with NetSquid made it a natural choice.
@@ IV.1 Trapped-ion modelling
We model the trapped-ion server using the NetSquid-based library NetSquid-TrappedIons [38]. This library was first used in [40] and is here introduced in more detail.
We model the decoherence of trapped-ion qubits over time using a collective Gaussian dephasing process [66], which can be rewritten as [40]
@@
The emission of entangled photons from an ion qubit is modelled in NetSquid-TrappedIons as the creation of a photon that has a polarisation degree of freedom that is maximally entangled with the state of the ion used in the emission, |$\ket{\psi}=(\ket{0H}+\ket{1V})/\sqrt{2}$, followed by the application of a single-qubit depolarising channel on the photon’s polarisation. This results in a Werner state of the form $\frac{3}{4}F|\psi\rangle\langle\psi|+\frac{4F-1}{3}\frac{\mathbb{1}}{4}$, where $F$ is the fidelity of the state with respect to the perfect state $|\psi\rangle$.
@@ IV.2 Client apparatus modelling
In simulating the client depicted in figure 1, the effect of the waveplates is given by the multiplication of the waveplate Jones matrices with the state vector [67]. The relative phase retardation induced between the fast axis and the slow axis is $\delta=\pi/2$ for a quarter waveplate and $\delta=\pi$ for a half waveplate. The fast axes of the waveplates also have an angle of $\xi$ radians with respect to the x-axis (which is along the plane of polarisation for linearly polarised light), determining the specific rotation that is implemented. However, errors in the setup can influence the retardation, giving a retardation deviation of $\Delta\delta$. It is also possible to have a deviation in the angle $\xi$ leading to $\Delta\xi$. With this, we can write the Jones matrices in general form [68] to include the errors as
@@
where $\delta^{\prime}=\delta+\Delta\delta$ (with $\delta=\pi/2$ for a QWP, and $\delta=\pi$ for a HWP) and $\xi^{\prime}=\xi+\Delta\xi$. Together, $\Delta\delta$ and $\Delta\xi$ lead to estimated waveplate error probability as given in table 1. The waveplates are implemented as a custom operation in NetSquid according to the Jones matrix.
Reference [43] gives an overview of general qubit rotations in terms of these fast axis settings. We can set the fast axes to correspond to a measurement in the |±$|\pm_{\theta}\rangle$-basis as
$\xi_{1}=0;$
$\xi_{2}=\theta/2;$
$\xi_{3}=\theta/4-3\pi/4.$
@@ IV.3 Minimisation methods
In our analysis, we use one target metric: the error tolerance of 25%, which is a bound on the fraction of test rounds that are allowed to fail while still being $\epsilon$-composably secure. The exact value of $\epsilon$ is not considered in this analysis, apart from that it can be made exponentially small by increasing the number of test rounds. In reference [32] the fraction of failed test rounds $w$ over the total number of test rounds $t$ is bounded by
@@
Here, $k$ is the principal colouring of the computation graph (which is the smallest number of ‘colours’ or labels one can give to the nodes in a graph such that no two neighbouring nodes have the same colour; see, for example, reference [69]) and $p$ is the inherent error probability of the bounded-error quantum computation. Assuming a computation for which $p=0$ we require the error probability to be below $1/2k$. The one-dimensional graph for single-qubit rotations used in this paper is two-colourable, which gives a maximum error tolerance of 25%.
@@
We look for absolute minimal hardware requirements by setting all but one parameter to perfect aside from fibre attenuation. For the coherence time, perfect means removing the collective dephasing noise model (section IV.1) from the ions in the trap. We then sweep over the imperfect parameter to find where the error probability due to this imperfection crosses the threshold.
@@
To find the crossing point, we do an initial global search with a small number of test rounds per point (to limit computation time) to find the approximate regime in which the error probability would pass the threshold. Once the region is located, the search is focused by taking the closest point above and below the threshold, halving the distance to their mean and running the simulation again for these points with a larger number of rounds. This process is repeated until the error-probability confidence interval of the points crosses the threshold of 25%. (This is similar to the bisection method used in root finding). The confidence intervals are determined by Hoeffding’s bound [70] as $\sqrt{\ln{(2/0.05)}/2t}$, with $t$ the number of test rounds. The focused search is executed with 70000 points in order to have a confidence level of 95% in an interval $\pm$ 0.005 for the error probability. The minimal requirements given in figure 4 are then extracted from the closest points ($(x_{1},y_{1}),(x_{2},y_{2})$) by a linear interpolation at the threshold ($y=0.25$) as $x=(y-y_{1})*(x_{2}-x_{1})/(y_{2}-y_{1})$. The error in these estimates is found by applying the same interpolation to the edges of the error probability confidence interval.
@@
Next, we find the set of minimal improvements. That is, from a given baseline (table 2), what parameters allows us to fulfill the constraint on the error probability with the least improvement? In order to quantify the cost of improving a parameter by a certain amount, we define hardware costs $H_{c}$ and a cost function $C$, which combines the hardware costs with our constraint on the error probability to give a single-objective minimisation problem as done in [39, 71, 72, 40]. We then employ a genetic algorithm using a workflow manager called YOTSE [73] to find the set of hardware parameters that minimises the cost function. This optimisation was run on SURF’s [^2] SURF is a collaborative organisation for IT in Dutch education and research. high-performance-computing cluster Snellius (Platinum 8360Y CPU @ 2.4GHz, maximum of 480 GB RAM) and on a workstation featuring an Intel Xeon Gold 6230 CPU @ 2.10GHz and 188 GB of DDR4 RAM, taking around 8000 core-hours per optimisation run of 20 generations.
@@
This is equivalent to summing over the improvement factor of each parameter. The probabilities of no-imperfection are defined for the optimisation parameters as in table 3.
@@
Server efficiency $\eta$ $\eta$
Coherence time $T_{c}$ $e^{-t^{2}/T_{c}^{2}}$
(Gate and entangled state) fidelity $F$ $\frac{1}{3}(4F-1)$
@@
For derivations and further explanation of the probabilities of no-imperfection, see Supplementary Note 6 of reference [40]. Note that the variable $t$ in the probability of no-imperfection of the coherence time, indicating the timescale over which qubits decohere, does not influence the hardware cost, as
$\frac{\ln\left(p_{NI}(b_{i})\right)}{\ln\left(p_{NI}(x_{i})\right)}=\frac{\ln\left(e^{-t^{2}/T_{c}^{2}}\right)}{\ln\left(\sqrt[k]{e^{-t^{2}/T_{c}^{2}}}\right)}=\frac{-t^{2}/T_{c}^{2}}{-t^{2}/kT_{c}^{2}}=k.$ (9)
@@
We now combine our requirement of having an error probability below 25% with the hardware cost to find the total cost of a set of parameters. We want the cost assigned to a set of parameters to be very high when the constraint is not met, and to be lower the closer parameter sets are to the baseline assuming that the constraint is met. A function to capture this behaviour, similar to what is used in references [39, 40, 71, 72], is
$C=w_{1}(1+(w/t-1/(2k))^{2})\Theta(w/t-1/(2k))+w_{2}Hc(x_{1},...,x_{N}),$ (10)
@@
where $\Theta(x)$ is the Heaviside step function and $w_{1}$ and $w_{2}$ are the weights of the objectives. We choose $w_{1}\gg w_{2}$ in order for the function to reflect that it is much more important to satisfy the error probability requirement than to minimise hardware cost, i.e., we do not care about the hardware cost as long as the requirement is not met.